OpenStudy (anonymous):

help plz (6x^3-1x^2+1)/(2x+1)

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

\(\huge \frac{6x^3-1x^2+1}{2x+1}=\frac{(2x+1)(3x^2-2x+1)}{2x+1}=3x^2-2x+1\)

OpenStudy (epoweritheta):

Friend use long division